OpenStudy (anonymous):

Two right triangles with the following dimensions are graphed with their hypotenuses on the same line: triangle X: Horizontal leg measures 2 units vertical leg measures 5 units triangle Y: Horizontal leg measures 7 units vertical leg measures 17.5 units Which other triangles could also have their hypotenuse on the same line? Choose all answers that are correct.

OpenStudy (anonymous):

A. horizontal leg measures 8 units, vertical leg measures 10 units B. horizontal leg measures 4 units, vertical leg measures 10 units C. horizontal leg measures 6 units, vertical leg measures 15 units D. horizontal leg measures 10 units, vertical leg measures 25 units @shinalcantara

OpenStudy (anonymous):

I think a and d

OpenStudy (shinalcantara):

you can use tangent with this:  you can use tangent in here to find the angle tan A = horizontal/vertical

OpenStudy (shinalcantara):

actually only 1 is not lying on the same line in here

OpenStudy (anonymous):

is it a and d

OpenStudy (shinalcantara):

try to evaluate y/x 5/2 = ? 17.5/7 = ? compare the resulting value to the values attained for the choices: A. 10/8 = ? B. 10/4 = ? C. 15/6 = ? D. 25/10 = ?

OpenStudy (anonymous):

a and d those r my anwers are they wrong

OpenStudy (shinalcantara):

let's solve it then. 5/2 = 2.5 17.5/7 = 2.5 so the slope of the line is 2.5 let's check the choices: A. 10/8 = 1.25 B. 10/4 = 2.5 C. 15/6 = 2.5 D. 25/10 = 2.5

OpenStudy (anonymous):

so b.c.d ??

OpenStudy (shinalcantara):

yep